OpenStudy (anonymous):

Can you show me the steps to solve -2(x-5)-7 please ?

OpenStudy (goldphenoix):

Do the distributive property. So what do you think you do first?

OpenStudy (anonymous):

I got -2x+10-7 ?

OpenStudy (goldphenoix):

Correct.

OpenStudy (goldphenoix):

What can you simplify next?

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay and then I subtracted -2x+10-10-7-10 ?I got -2x=-17

OpenStudy (anonymous):

to finish it u must add like terms @Amanda01

OpenStudy (anonymous):

that one is corrct :)

OpenStudy (goldphenoix):

What? Wait. Why you subtract?

OpenStudy (anonymous):

wats with the last one @Amanda01

OpenStudy (anonymous):

I did it wrong :/

OpenStudy (goldphenoix):

So your equation looks like: -2x+10-7= Correct? Do you see any terms that we can combine?

OpenStudy (anonymous):

the 10 and the 7 ?

OpenStudy (goldphenoix):

Correct!

OpenStudy (goldphenoix):

What would your answer be?

OpenStudy (anonymous):

so 10-7 which is 3 ?

OpenStudy (goldphenoix):

Yes. So your final answer would be: -2x+3. Unless your equation in the first place was: -2(x-5)-7=0. What grade are you in?

OpenStudy (goldphenoix):

Oh wait a minute. You want to solve for x?

OpenStudy (anonymous):

your gonna laugh when I tell you but im actually supposed to be graduating.. my math is what is holding me back . Im HORRIBLE with math ..

OpenStudy (goldphenoix):

It's okay. Are you in high school or middle school?

OpenStudy (anonymous):

im just got out of my senior year in high school

OpenStudy (goldphenoix):

Hmm. Since you want to solve for x: The equation should look like: -2x + 10 + -7 = 0 What can we simplify in the left side?

OpenStudy (anonymous):

it was supposed to say -2(x-5)=7

OpenStudy (goldphenoix):

Yes. But you already distribute it. :) So our equation we got so far was -2x + 10 + -7 Since we want to find x, you put =0. So what do we do next?

OpenStudy (anonymous):

-10 ?

OpenStudy (goldphenoix):

-2x + 10 + -7 = 0 Do you see any terms in the left side that we can combine?

OpenStudy (anonymous):

10 and 7

OpenStudy (goldphenoix):

Correct! So what is 10 + - 7 give you?

OpenStudy (anonymous):

3

OpenStudy (goldphenoix):

Correct!! Now our equation looks like: -2x + 3 = 0 What do you think we do next?

OpenStudy (anonymous):

subtract 3 from both sides ?

OpenStudy (goldphenoix):

Yep. What does your equation looks like now?

OpenStudy (anonymous):

-2x=-3

OpenStudy (goldphenoix):

Correct! What would the final step be?

OpenStudy (anonymous):

divide ?

OpenStudy (goldphenoix):

Yes. Divide what number to both side?

OpenStudy (anonymous):

-2

OpenStudy (goldphenoix):

Correct!!! So what does x give you?

OpenStudy (anonymous):

x=-1/2

OpenStudy (goldphenoix):

Nope. What is 3 divided by 2?

OpenStudy (anonymous):

1.5

OpenStudy (goldphenoix):

Correct! Therefore, x = 1.5. Since a negative divided by a negative is a positive. :) Do you have any question that you want to ask?

OpenStudy (anonymous):

I did -3 divided by -2 in my cal and is gave me -1.5 but idk .. lol . You are such an awesome help. Thank you so much

OpenStudy (goldphenoix):

Are you sure? -3 divided by -2 will give you the same answer as 3 divided by 2. :)

OpenStudy (goldphenoix):

Yeah, no problem. I hope this help! :)

OpenStudy (anonymous):

It did . thanks for your time ! (:

OpenStudy (goldphenoix):

Yes.! I hope you are able to pass in math. :) If you need any help, ask me. :) (Also, medal please. >_> I'm greedy for medal. c:)

OpenStudy (anonymous):

How do I medal ?

OpenStudy (goldphenoix):

Do you see best response or something? ------------------------------->

OpenStudy (anonymous):

oh , yup (: